OpenStudy (anonymous):

if each of the four letters in the sum below represents a different digit, which of the following cannot be a value of A? AB + BA ------ CDC 1. 6 2. 5 3. 4 4. 3 5. 2

OpenStudy (anonymous):

Answer 5.

OpenStudy (saifoo.khan):

\[\Huge \color\violet 5\]

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

the answer isnt 5, you can have 56+ 65 = 121

OpenStudy (anonymous):

No, as in choice 5.

OpenStudy (anonymous):

The number AB can be represented as 10A+B BA = 10B+A Their sum is 10(A+B) + (A+B) = 11(A+B) The only way for 11(A+B) to have an answer like CDC is is A + B = 11

OpenStudy (anonymous):

So you try the answers out. A = 6 means B = 5, 65+56 = 121 thats fine, they are all different digits. You test them all, they will all be fine, except for A = 2 A = 2 means B = 9, we get 29+92 = 121, but D cant equal 2 because A is already 2.